40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  CTN Lab elog  Not logged in ELOG logo
Message ID: 214     Entry time: Wed Jul 14 19:07:17 2010
Author: Megan 
Type: DailyProgress 
Category: VCO 
Subject: VCO Update 

I replace R3 and R7 with thin film resistors (according to LISO they have the greatest contribution of the resistors at lower frequencies) and re-measured the electronic noise of the driver. It was almost identical to the noise I previously measured, and so for now am not going to replace the other resistors because it does not seem to have a large effect.

I have been playing around with the LISO model more and will hopefully have an easy-to-use version to upload soon. I'm going to work with more of the data that I have to see what it tells us and will post various graphs tomorrow.

ELOG V3.1.3-